OpenStudy (anonymous):

convert 1.8 radians into degrees

OpenStudy (richyw):

\[\pi \text{ rad}=180^{\circ}\]\[1.8\text{ rad}\times\frac{180^{\circ}}{\pi\text{ rad}}=324^{\circ}\]

OpenStudy (doc.brown):

\[\tau=360^o\]\[6.283185...rad=360^o\]Through ratio, using cross multiplication: \[\frac{1.8}{\tau}=\frac{x}{360}\]Through conversion:\[1.8rad\times\frac{360^o}{6.28rad}=\]

OpenStudy (anonymous):

103.13?

OpenStudy (doc.brown):

Does the question ask for a certain number of decimal places?

OpenStudy (doc.brown):

@richyw That answer is incorrect.

OpenStudy (anonymous):

i got it right, its 103.13 thank you!

OpenStudy (doc.brown):

If this was science, and not math, 1.8 has 2 significant figures which means your answer can only have 2 significant figures.\[1.0\times 10^2deg\]103.13 is correct enough for math class.

OpenStudy (richyw):

oops. forgot to divide by pi! good catch. I was in a hurry to leave and fix my truck. the method was correct though